OpenStudy (anonymous):

3+y/4=-y/8 solve the proportion

OpenStudy (anonymous):

\[(12+y)/4 = -y/8\]

OpenStudy (anonymous):

\[24+2y=-y\]

OpenStudy (anonymous):

\[y=-8\]

OpenStudy (anonymous):

Is this okay?